# Too Many Frames, Not All Useful: Efficient Strategies for Long-Form Video QA

### v4

@@ Appendix A.2 Algorithms in Detail
1:Require: ResNet-18 [11] pretrained on imagenet dataset $f$, frame list $\textbf{List}_{frame}$, image index list $\textbf{List}_{index}\in\{1,\dots,N\}$, minimum number of list length $\psi$, temperature $\lambda$, number of sample $\tau$, function to find index of x in list w index(x, w), and function to sort list sort(List)
@@
7:   $\textbf{List}_{dist}\leftarrow\frac{\sum_{y}\sum_{x}{\sqrt{(\mathbf{F_{i}}-\textbf{List}_{feat})^{2}}}}{x\times y}$
@@
13:   $i\leftarrow\textbf{List}_{index}$.pop(00)
14:   $\textbf{p}^{i}\leftarrow\texttt{softmax}(\textbf{M}^{i}_{dist})$
15:   $\mu_{\mathbf{\textbf{p}^{i}}},\sigma_{\mathbf{\textbf{p}^{i}}}\leftarrow\texttt{mean}(\textbf{p}^{i}),\texttt{std}(\textbf{p}^{i})$
16:   $\beta\leftarrow\mu_{\mathbf{\textbf{p}^{i}}}-\sigma_{\mathbf{\textbf{p}^{i}}}\sum_{i=0}{e^{1-i/\lambda}}$
@@
23:      $\delta$leftarrow $\gamma~{}th$ value in $\textbf{List}_{index}$
@@
26:   end for
27:   $\textbf{List}_{\delta}$.insert($i$)
28:   $\textbf{List}_{sample}$leftarrow sample $\tau$ items from $\textbf{List}_{\delta}$
29:   $\textbf{sort}(\textbf{List}_{sample})$
@@
1:Require: keyword set $\mathbf{K}$, image set $\mathbf{I}$, total length of selected image set $L$, function to calculate similarity matrix $\textbf{sim}(\mathbf{K},\mathbf{I})$, function to sort similarity matrix and return indices $\textbf{sort}(\mathbf{S})$
2:$\mathbf{S}\leftarrow\textbf{sim}(\mathbf{K},\mathbf{I})$
3:$\mathbf{S}_{sorted},\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
@@
5:Initialize $\mathbf{I}_{selected}$ as an empty set
@@
20:end while
21:return $\mathbf{P}_{best}$
@@
2:$\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
@@
4:$\mathbf{idx}_{temporal}\leftarrow\textbf{temporal\_order}(\mathbf{idx}_{low\_sim})$
@@
6:$\mathbf{I}_{ordered},\mathbf{K}_{ordered}\leftarrow\mathbf{I}[\mathbf{idx}_{final}],\mathbf{K}[\mathbf{idx}_{final}]$
@@
10:end for
11:return $\mathbf{I}_{selected}$

### v5

@@ Appendix A.3 Algorithms in Detail
1
21:Require: ResNet-18 He et al. (2016b) pretrained on imagenet dataset $f$, frame list $\textbf{List}_{frame}$, image index list $\textbf{List}_{index}\in\{1,\dots,N\}$, minimum number of list length $\psi$, temperature $\lambda$, number of sample $\tau$, function to find index of x in list w index(x, w), and function to sort list sort(List)
3
@@
67:   $\textbf{List}_{dist}\leftarrow\frac{\sum_{y}\sum_{x}{\sqrt{(\mathbf{F_{i}}-\textbf{List}_{feat})^{2}}}}{x\times y}$
@@
813:   $i\leftarrow\textbf{List}_{index}$.pop(00)
914:   $\textbf{p}^{i}\leftarrow\texttt{softmax}(\textbf{M}^{i}_{dist})$
1015:   $\mu_{\mathbf{\textbf{p}^{i}}},\sigma_{\mathbf{\textbf{p}^{i}}}\leftarrow\texttt{mean}(\textbf{p}^{i}),\texttt{std}(\textbf{p}^{i})$
1116:   $\beta\leftarrow\mu_{\mathbf{\textbf{p}^{i}}}-\sigma_{\mathbf{\textbf{p}^{i}}}\sum_{i=0}{e^{1-i/\lambda}}$
@@
1323:      $\delta$leftarrow $\gamma~{}th$ value in $\textbf{List}_{index}$
@@
1426:   end for
1527:   $\textbf{List}_{\delta}$.insert($i$)
1628:   $\textbf{List}_{sample}$leftarrow sample $\tau$ items from $\textbf{List}_{\delta}$
1729:   $\textbf{sort}(\textbf{List}_{sample})$
@@
1
21:Require: keyword set $\mathbf{K}$, image set $\mathbf{I}$, total length of selected image set $L$, function to calculate similarity matrix $\textbf{sim}(\mathbf{K},\mathbf{I})$, function to sort similarity matrix and return indices $\textbf{sort}(\mathbf{S})$
3
42:$\mathbf{S}\leftarrow\textbf{sim}(\mathbf{K},\mathbf{I})$
53:$\mathbf{S}_{sorted},\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
@@
65:Initialize $\mathbf{I}_{selected}$ as an empty set
@@
720:end while
821:return $\mathbf{P}_{best}$
@@
3
42:$\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
@@
64:$\mathbf{idx}_{temporal}\leftarrow\textbf{temporal\_order}(\mathbf{idx}_{low\_sim})$
@@
86:$\mathbf{I}_{ordered},\mathbf{K}_{ordered}\leftarrow\mathbf{I}[\mathbf{idx}_{final}],\mathbf{K}[\mathbf{idx}_{final}]$
@@
1010:end for
$1$:return $\mathbf{I}_{selected}$
@@ Appendix A.4 Prompting: Fine Keyframe Detector
Appendix A.5 Comparison with Other Keyframe Selection Methods
We aim to highlight the main advantage of the Hierarchical Keyframe Selector over other existing keyframe selection methods. Models like VideoAgent, VideoTree, and TraveLER provide useful comparisons, as they utilize keyframe selection mechanism with similar or different scale of frames. VideoAgent and TraveLER rely on uniform frame selection in the first iteration without analyzing the entire video even though they perform non-uniform sampling in the next iterations. They identify important segments based solely on these initial frames and the LLM’s response, which can be problematic if the initial uniformly selected frames are not representative of the entire video or if the LLM misinterprets the captions and prompts. In such cases, the LLM might incorrectly identify segments for further analysis. If the LLM fails to pinpoint the correct segment initially, the entire process can break down because subsequent frames will be similar to the first set, leading the LLM to continuously select frames within or near the initial segment. Additionally, for videos that are as challenging or more difficult than EgoSchema in terms of temporal complexity and activities, existing keyframe selection models such as VideoAgent, VideoTree, and TraveLER may require numerous iterations by running heavy visual/language models to finalize keyframes selection. This results in higher computational and latency costs, as it necessitates numerous runs of resource-intensive VLM and LLM models.
In contrast, our method analyzes the entire video with high frame rates using a lightweight ResNet-18 (He et al., 2016a) and segments the video non-uniformly based on scene continuity. We then select several frames in each segment by measuring feature similarity between frame features and keywords using the CLIP-B/16 (0.12B) (Ranasinghe et al., 2023) which is lighted than VideoAgent’s EVA-CLIP-8Bplus (8B). By reviewing the entire video and non-uniformly selecting keyframes based on scene continuity and similarity scores, these keyframes accurately represent the question-based important frames distribution in the entire video. Furthermore, we use VLM for a fine-grained selection of keyframes, improving keyframe selection when CLIP-B/16 struggles to understand detailed atomic activities in the frames. By hierarchically segmenting the video with different modules, the resulting segments and keyframes are more reliable than those from VideoAgent. Even with more challenging videos, our process only needs to go through the video once to collect keyframes, maintaining computational efficiency.
Figure A.5 visualizes the differences of the keyframe selection mechanism bewtween LVNet and VideoAgent. On the left, LVNet begins with uniformly sampled frames and filters them through multiple stages, resulting in a non-uniform distribution of frames over time. First, the temporal scene clustering (TSC) selects somes frames that represent temporally distinct activities. Next, the coarse keyframe detector (CKD) targets frames most relevant to the question. Finally, the fine keyframe detector (FKD) further refines this selection to ensure the keyframes accurately capture the activity in question. As a result, LVNet produces 12 frames, with 8 of them (67%) directly depicting "usage of phones," which is the correct answer and leads the model to select the right option. On the right, VideoAgent also starts with the uniform frames but relies on a LLM to request additional frames. Since the initial frames do not capture enough relevant content, the LLM again selects frames uniformly, adding more irrelevant samples that lack the crucial information about "usage of phones." As a result, VideoAgent ultimately selects the wrong option.